A nearsighted person has a near point of 14 cm and a far point of 40 cm.

What power corrective lens is needed for her to have clear distant vision?

With this corrective lens in place, what is her new near point?
Answer #1

(a) for near sight :

P = -100/PR ----> PR = the far point

= -100/40 = - 2.5 D

(b) f = 100/P = 100/-2.5 = -40 cm

di = -14 cm

--> do = di * f/(di - f)

= (-14 * -40)/(-14 - (-40))

= 21.5 cm